# Quantum Dynamics Predicts Coherent Oscillatory Behavior in the Early-times of a Photoisomerization Reaction

### v1

@@ 1 Introduction
Characterizing the quantum dynamics of a molecular system in the excited states is a challenging task, even when reduced dimensionality models are employed to restrict the available configurational space to well-chosen representative “reaction coordinates”. Often, the selection of such coordinates is guided by the behavior of the molecule in the Franck-Condon region and by normal modes analysis1, 2, 3, 4, but clearly (and not surprisingly), these modes fail capturing salient dynamic features when these involves large amplitude motions. This is indeed the case of photoisomerization reactions, which are phenomena abundantly explored in the literature for their multiple applications, ranging from photochromism5 to photoswitches6, 7 and for their role in photobiology8, 9.
@@ 2 Presentation of the model and computational details
The quantum dynamics (QD) of photoexcited cis-PSB3 is simulated here for 200 fs using the Quantics code14 in the electronic diabatic basis (we recall that the model itself accounts for two electronic diabatic states of covalent and charge-transfer character11). The Short Iterative Lanczos method is used, increasing the size of the basis until convergence is ensured. In this work, we only adopted the approximate kinetic energy operator of Ref.[13] with a constant and diagonal metric tensor, whose components are a.u., a.u., a.u. As in Ref.[13], the initial vibrational wavepacket (WP) is chosen to be a three-dimensional Gaussian centered at with zero initial momentum and width a.u. along the three vibrational directions.
@@
Figure 3 shows the adiabatic populations of S0 (lower panels) and S1 (upper panels) as functions of time, resolved in the cis (left panels) and trans (right panels) isomers by integrating the nuclear probability densities in S0 and in S1 over all the BLA and HOOP domains but in the interval for the cis isomer and outside of this interval for the trans isomer. In Figure 3, we report for comparison the results of Ref.[13] (light-green curves), that were obtained using grid points for BLA, Tors and HOOP, respectively. We note here that, while the behavior of the population in S1 is qualitatively fairly stable with respect to the increase of the grid size, the S0 populations of both isomers change qualitatively, and approach convergence only after a certain threshold (in the order magenta, light-green, orange, and dark-green, the curves correspond to 94, 136, 162, 192 and show a somehow non-monotonic convergence to the thick gray lines). In particular, the results suggest that when QD along Tors is not correctly described, the WP moving in the trans region is prevented from repopulating the cis isomer, thus missing the clear oscillations observed in the lower left panel of Figure 3.
@@ 3 Quantum and quantum-classical dynamics
Figure 4 shows, similarly to Figure 3 above, the evolution of (upper left panel), (upper right panel), (lower left panel), (lower right panel). The QD results of the present work are shown in black, along with the previous results of Ref.[13] in light-green. Quantum-classical results have been obtained using the trajectory surface hopping method (TSH),16 also including energy-based decoherence (TSH-EDC) corrections,17 the multi-trajectory Ehrenfest method, and coupled-trajectory mixed quantum-classical (CT-MQC) algorithm.18 In Figure 4, red and orange lines indicate TSH and TSH-EDC, respectively, whereas purple lines correspond to Ehrenfest and blue lines to CT-MQC. Trajectory-based dynamics was performed in Ref.[13] using the G-CTMQC code19 interfaced with the QuantumModelLib potential library20 where the PSB3 model is implemented.
@@
The largest deviation of our QD results from the original QD results of Ref.[13] is observed for the Tors kinetic energy. The new findings suggest that the differences in the cis and trans populations stem from an underestimation of the Tors kinetic energy in the original simulations, due to the small size of the basis set used along this mode. Note that within the first 50 fs of dynamics all predictions are similar and they all show that the amplitude of the oscillations in the BLA kinetic energy is damped while the WP spreads toward the conical intersections at .
After 50 fs, the quantum-classical methods predict that the Tors motion acquires kinetic energy which seems largely overestimated when compared to the original QD results, with a maximum difference reached at 80 fs delay. Instead, the converged QD results agree very well with the quantum-classical ones. The larger kinetic energy gained by Tors according to the new QD results is compensated by a smaller kinetic energy along the BLA and HOOP. A more detailed analysis suggests that, as observed above for the populations and the quantum yield, TSH and TSH-EDC predictions are in very good agreement with the QD reference as far as Tors is concerned, whereas, at least after 100 fs CT-MQC gives a more accurate prediction for the HOOP kinetic energy. In addition, the loss in BLA kinetic energy and the damping of its oscillations predicted by the new QD is fairly well captured by the quantum-classical methods. Such features are likely due to a more effective decoherence in the BLA direction, due to the larger WP spreading along Tors and HOOP coordinates (as will be discussed below). As discussed above, the large kinetic energy available along the Tors coordinate is responsible for the transcis back-reaction in S0 and, therefore, for the oscillatory behavior of the ground-state trans population and of the quantum yield.
@@ 4 Conclusions
Photoisomerizations play a key role in many chemical and biological processes and, when occurring in the ultrafast subpicosecond regime, their mechanism is usually characterized by the existence of conical intersections and nonadiabatic couplings between electronic states. They are, therefore, an intrinsically quantum-mechanical process whose natural description is obtained through quantum dynamics. However, quantum dynamics simulations are often time consuming or even prohibitive, making trajectory-based methods a very attractive alternative. Additionally, the application of trajectory-based methods, once properly validated, may offer the possibility to unveil details of the reaction mechanism that are not easily accessible in a quantum dynamics framework. This was exactly the case of the analysis of the photoisomerization in cis-PSB3,13 where the trajectories predict, even in a dramatically reduced model chromophore, the existence of a phase relationship between Tors and HOOP velocities whose signs at the time of the transition from S1 to S0 determine the success of the isomerization in a scotopic biological photoreceptor 21, 8, 22, 23. These considerations highlight the need to provide the community with reliable reference quantum dynamics data for the model of Refs.[11, 13]. The present work demonstrates an excellent agreement between quantum dynamics and surface-hopping dynamics, an observation that has been revisited here with respect to Ref.[13].
@@
Concerning the information on the photoisomerization reaction extracted from the quantum dynamics, the results presented in this work suggest the appearance of a first coherent oscillatory behavior in the early-time relaxation towards the electronic ground state forming the cis isomer. The initial appearance of population of the cis isomer in S0 is about 50 fs and the oscillations in the population agree with the oscillations of the kinetic energy along the HOOP mode, thus, these results seem to suggest that HOOP plays a key role in guiding the relaxation after photoexcitation and the isomerization reaction, as was observed for the same model from the analysis of representative classical trajectories in Ref.[13] and also previously when reporting on the multi-scale modelling of the dynamics of the entire biological photoreceptor 21, 8, 22, 23.
As stressed also in Ref.[13], the model does not take into consideration either additional (internal) vibrational modes or the environment, and due to the reactant planarity, population transfer from S1 to S0 takes place symmetrically along the clockwise and counterclockwise directions Tors 0 and Tors 0. Furthermore, the energy pumped into the system via the initial electronic excitation cannot be dissipated, causing the nuclear wavepacket to remain in S1, as it is evident in all the numerical results present in this work.
Our analysis of the quantum dynamics suggests that the majority of the population of the initially populated state decays initially to S0 in the trans isomer, reaching a maximum at 80 fs, after which a second coherent oscillatory feature drives a cis population increase in S0 up to 125 fs, which is accompanied by a decrease of the trans isomer in S0. This implies, as discussed above, the transcis back-reaction in the ground state. After 130 fs, the cis population in S0 decays in favor of the trans population in S0. Specifically, these results are evidence of the occurrence of a trans/cis interconversion in the ground state, since at times larger than fs the cis and trans populations on S1 undergo only mild changes, respectively decreasing and increasing by 0.05 up to 200 fs.
@@ Acknowledgements
MA, MG and FS gratefully acknowledge financial support from the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Chemical Sciences, Geosciences and Biosciences Division under award no. DE-SC0022225. FS and MG acknowledge support from the CRESCENDO project, PRIN: PROGETTI DI RICERCA DI RILEVANTE INTERESSE NAZIONALE-Bando 2022, Prot. 2022HL9PRP. FA acknowledges financial support from the French Agence Nationale de la Recherche under the grants No. ANR-20-CE29-0014 (ANR Q-DeLight project) and No. ANR-23-ERCC-0002 (ANR STROM project).

### v2

@@ 1 Introduction
Characterizing the quantum dynamics of a molecular system in the excited states is a challenging task, even when reduced dimensionality models are employed to restrict the available configurational space to well-chosen representative “reaction coordinates”. Often, the selection of such coordinates is guided by the behavior of the molecule in the Franck-Condon region and by normal modes analysis1, 2, 3, 4, but clearly (and not surprisingly), such modes fail capturing salient dynamic features when these involve large amplitude motions. This is indeed the case of photoisomerization reactions, which are phenomena abundantly explored in the literature for their multiple applications, ranging from photochromism5 to photoswitches6, 7 and for their role in photobiology8, 9.
@@ 2 Presentation of the model and computational details
The quantum dynamics (QD) of photoexcited cis-PSB3 is simulated here for 200 fs using the Quantics code14 in the electronic diabatic basis (we recall that the model itself accounts for two electronic diabatic states of covalent and charge-transfer character11). The Short Iterative Lanczos (SIL) method is used, increasing the size of the basis until convergence is ensured. In this work, we only adopted the approximate kinetic energy operator of Ref.[13] with a constant and diagonal metric tensor, whose components are a.u., a.u., a.u. As in Ref.[13], the initial vibrational wavepacket (WP) is chosen to be a three-dimensional Gaussian centered at with zero initial momentum and width a.u. along the three vibrational directions.
@@
Figure 3 shows the adiabatic populations of S0 (lower panels) and S1 (upper panels) as functions of time, resolved in the cis (left panels) and trans (right panels) isomers by integrating the nuclear probability densities in S0 and in S1 over all the BLA and HOOP domains but in the interval for the cis isomer and outside of this interval for the trans isomer. In Figure 3, we report for comparison the results of Ref.[13] (light-green curves), that were obtained using grid points for BLA, Tors and HOOP, respectively. We note here that, while the behavior of the population in S1 is qualitatively fairly stable with respect to the increase of the grid size, the S0 populations of both isomers change qualitatively, and approach convergence only after a certain threshold (in the order magenta, light-green, orange, and dark-green, the curves correspond to 94, 136, 162, 192 and show a somehow non-monotonic convergence to the thick gray lines). In particular, the results suggest that when QD along Tors is not correctly described, the WP moving in the trans region is prevented from repopulating the cis isomer, thus missing the clear oscillations of the S0 trans population with a period of 100 fs observed in the lower right panel of Figure 3. Similarly, the fast oscillations of the S0 cis population (lower left panel) are non correctly reproduced.
The need for such a large primitive basis set suggests that the adoption of more efficient propagation methods than SIL, like Multiconfigurational Time Dependent Hartree (MCTDH),16, 17 might reveal particularly beneficial. Testing the performance of MCTDH in this three-dimensional photoisomerization model is also important since its better scalability with the number of coordinates (especially in the multilayer variant 18, 19) makes it possibly applicable to study the QD of photoisomerizations involving a larger number of coordinates.
We performed extensive numerical tests, which are detailed in Appendix A. There, we show that convergence to the exact numerical results is possible both in the single-set and in the multi-set formalism,17 the latter being more efficient in terms of computational time. The better performance of the multi-set ansatz can be rationalized due to the large differences between the ground and excited potential energy surfaces, which make specific single-particle functions (SPFs) more accurate in describing the WP. Interestingly, we demonstrate that achieving convergence requires a remarkably large number of SPFs, especially for the TORS coordinate in the ground state, whereas capturing the dynamics along BLA and HOOP is less demanding. Specifically, we used 160 and 110 SPFs along the Tors mode in the ground and in the excited state, respectively. An insufficient number of SPFs leads to notable inaccuracies in both diabatic and adiabatic populations after 100 fs and hinders the proper exchange of trans/cis population observed in fully converged simulations (discussed below). This effect is also reflected in the computed quantum yield, emphasizing the need for careful convergence testing when using MCTDH for strongly anharmonic systems, especially if a large amount of kinetic energy accumulates in few degrees of freedom.
@@ 3 Quantum and quantum-classical dynamics
Figure 4 shows, similarly to Figure 3 above, the evolution of (upper left panel), (upper right panel), (lower left panel), (lower right panel). The QD results of the present work are shown in black, along with the previous results of Ref.[13] in light-green. Quantum-classical results have been obtained using the trajectory surface hopping method (TSH),20 also including energy-based decoherence (TSH-EDC) corrections,21 the multi-trajectory Ehrenfest method, and coupled-trajectory mixed quantum-classical (CT-MQC) algorithm.22 In Figure 4, red and orange lines indicate TSH and TSH-EDC, respectively, whereas purple lines correspond to Ehrenfest and blue lines to CT-MQC. Trajectory-based dynamics was performed in Ref.[13] using the G-CTMQC code23 interfaced with the QuantumModelLib potential library24 where the PSB3 model is implemented.
@@
The largest deviation of our QD results from the original QD results of Ref.[13] is observed for the Tors kinetic energy. The new findings suggest that the differences in the cis and trans populations stem from an underestimation of the Tors kinetic energy in the original simulations, due to the small size of the basis set used along this mode. Note that within the first 50 fs of dynamics all predictions are similar and they all show that the amplitude of the oscillations in the BLA kinetic energy is damped while the WP spreads toward the conical intersections at . After 50 fs, the quantum-classical methods predict that the Tors motion acquires kinetic energy which seems largely overestimated when compared to the original QD results, with a maximum difference reached at 80 fs delay. Instead, the converged QD results agree very well with the quantum-classical ones. The larger kinetic energy gained by Tors according to the new QD results is compensated by a smaller kinetic energy along the BLA and HOOP. A more detailed analysis suggests that, as observed above for the populations and the quantum yield, TSH and TSH-EDC predictions are in very good agreement with the QD reference as far as Tors is concerned, whereas, at least after 100 fs CT-MQC gives a more accurate prediction for the HOOP kinetic energy. In addition, the loss in BLA kinetic energy and the damping of its oscillations predicted by the new QD is fairly well captured by the quantum-classical methods. Such features are likely due to a more effective decoherence in the BLA direction, due to the larger WP spreading along Tors and HOOP coordinates (as will be discussed below). As discussed above, the large kinetic energy available along the Tors coordinate is responsible for the transcis back-reaction in S0 and, therefore, for the oscillatory behavior of the ground-state trans population and of the quantum yield.
@@ 4 Conclusions
Photoisomerizations play a key role in many chemical and biological processes and, when occurring in the ultrafast subpicosecond regime, their mechanism is usually characterized by the existence of conical intersections and nonadiabatic couplings between electronic states. They are, therefore, an intrinsically quantum-mechanical process whose natural description is obtained through quantum dynamics. However, quantum dynamics simulations are often time consuming or even prohibitive, making trajectory-based methods a very attractive alternative. Additionally, the application of trajectory-based methods, once properly validated, may offer the possibility to unveil details of the reaction mechanism that are not easily accessible in a quantum dynamics framework. This was exactly the case of the analysis of the photoisomerization in cis-PSB3,13 where the trajectories predict, even in a dramatically reduced model chromophore, the existence of a phase relationship between Tors and HOOP velocities whose signs at the time of the transition from S1 to S0 determine the success of the isomerization in a scotopic biological photoreceptor 25, 8, 26, 27. These considerations highlight the need to provide the community with reliable reference quantum dynamics data for the model of Refs.[11, 13]. The present work demonstrates an excellent agreement between quantum dynamics and surface-hopping dynamics, an observation that has been revisited here with respect to Ref.[13].
@@
Concerning the information on the photoisomerization reaction extracted from the quantum dynamics, the results presented in this work suggest the appearance of a first coherent oscillatory behavior in the early-time relaxation towards the electronic ground state forming the cis isomer. The initial appearance of population of the cis isomer in S0 is about 50 fs and the oscillations in the population agree with the oscillations of the kinetic energy along the HOOP mode, thus, these results seem to suggest that HOOP plays a key role in guiding the relaxation after photoexcitation and the isomerization reaction, as was observed for the same model from the analysis of representative classical trajectories in Ref.[13] and also previously when reporting on the multi-scale modelling of the dynamics of the entire biological photoreceptor 25, 8, 26, 27.
As stressed also in Ref.[13], the model does not take into consideration either additional (internal) vibrational modes or the environment, and due to the reactant planarity, population transfer from S1 to S0 takes place symmetrically along the clockwise and counterclockwise directions Tors 0 and Tors 0. Furthermore, the energy pumped into the system via the initial electronic excitation cannot be dissipated, causing the nuclear wavepacket to remain in S1, as it is evident in all the numerical results present in this work. Our analysis of the quantum dynamics suggests that the majority of the population of the initially populated state decays initially to S0 in the trans isomer, reaching a maximum at 80 fs, after which a second coherent oscillatory feature drives a cis population increase in S0 up to 125 fs, which is accompanied by a decrease of the trans isomer in S0. This implies, as discussed above, the transcis back-reaction in the ground state. After 130 fs, the cis population in S0 decays in favor of the trans population in S0. Specifically, these results are evidence of the occurrence of a trans/cis interconversion in the ground state, since at times larger than fs the cis and trans populations on S1 undergo only mild changes, respectively decreasing and increasing by 0.05 up to 200 fs.
@@ Acknowledgements
MA, MG and FS gratefully acknowledge financial support from the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Chemical Sciences, Geosciences and Biosciences Division under award no. DE-SC0022225. FS and MG acknowledge support from the CRESCENDO project, PRIN: PROGETTI DI RICERCA DI RILEVANTE INTERESSE NAZIONALE-Bando 2022, Prot. 2022HL9PRP. M.O. is grateful for partial support provided by EU funding within the MUR PNRR “National Center for Gene Therapy and Drugs based on RNA Technology” (Project no. CN00000041 CN3 RNA) - Spoke 6. M.O. is also grateful to the NSF CHE-SDM A for Grant No. 2102619 and to the MUR for a PRIN 2022 grant No. 2022K3AY2K. FA acknowledges financial support from the French Agence Nationale de la Recherche under the grants No. ANR-20-CE29-0014 (ANR Q-DeLight project) and No. ANR-23-ERCC-0002 (ANR STROM project).
Appendix A MCTDH simulations
Although the numerically exact solution of the time-dependent Schrödinger equation on a full direct-product grid offers a highly accurate representation of the wavepacket (WP), its dimensionality, and consequently the computational cost, increases exponentially with the number of degrees of freedom. This exponential scaling limits the applicability of exact quantum propagation to relatively small systems, typically with fewer than 5–6 degrees of freedom. Thanks to the impressive methodological progress of Multi-Configurational Time-Dependent Hartree (MCTDH)17 and its multi-layer (ML) extension (ML-MCTDH),18, 19 handling systems with a large number of degrees of freedom (typically up to 20-30 degrees of freedom with MCTDH and up to hundreds with its ML extension) is now feasible. MCTDH exploits a variationally optimized, time-dependent basis to represent the WP. In practice, the WP is represented as the sum of products of time-dependent single-particle functions (SPFs). Therefore, the accuracy of the dynamics in MCTDH depends not only on the number of primitive basis functions used for each degree of freedom, but also, and strongly, on the number of SPFs.
In this Appendix, we systematically assess the convergence of the quantum dynamics simulation performed with MCTDH on the three-dimensional potential model describing the photoisomerization of PSB3. In Figure 8, we report the time evolution of the population of the diabatic (left) and of the adiabatic (right) states. As indicated in the figure, the various curves refer to different combinations of the numbers of SPFs adopted for BLA, Tors, and HOOP coordinates. We employed the multi-set formalism, which allows assigning different sets of SPFs to each vibrational mode on different electronic states. For comparison, we also performed the dynamics with the single-set formalism for a specific case, indicating that the multi-set formalism yields results equivalent to the single-set approach when using the same number of SPFs for both states. Moreover, for MCTDH simulations, we slightly increased the number of grid points for BLA and HOOP to provide more flexibility in assigning different numbers of SPFs. As shown in Figure 8, achieving convergence requires a large number of SPFs (up to 160) for the Tors mode in the ground state, whereas approximately 110 SPFs are sufficient in the excited state. This behavior is expected, given the significant amount of kinetic energy acquired by the Tors mode in the ground state. It is also evident that the dynamics is not particularly sensitive to the number of SPFs assigned to the HOOP and BLA modes, as seen by comparing the light-green solid line with the orange and yellow lines, where the number of SPFs for BLA and HOOP in the excited state was progressively reduced. A significant deviation from the reference results (in black), consistently observed after 100 fs, occurs when a reduced number of SPFs is used for Tors, particularly in the ground state.
Figure 9 shows, similarly to Figure 3, the time evolution of the populations (upper left panel), (upper right panel), (lower left panel) and (lower right panel) calculated using MCTDH with different numbers of SPFs assigned to each mode on different electronic states. Interestingly, the S1 population for both isomers appears to be more sensitive to the basis size (i.e., the number of SPFs) than to the primitive grid size, as evidenced by the significant deviation from the reference observed after 100 fs. A similar trend is observed for the S0 population, with the cis isomer showing a strong dependence on the number of SPFs in both the S0 and S1 states. Specifically, an insufficient number of SPFs appears to hinder the cis-to-trans population transfer in S0 observed in the fully converged quantum dynamics simulation after 130 fs. This effect is also reflected in the time-dependent quantum yields shown in Figure 10, where the overestimated cis population in S0 after 130 fs, caused by insufficient convergence, leads to a reduced quantum yield at long times.